Operations Report – January 8th

Crew 186 Operations Report 08JAN2018

SOL: 08

Name of person filing report: M. Grande

Non-nominal systems: Generator power system

Notes on non-nominal systems: Generator system limping along with a now-nominal routine. Early morning failures still attributed to the cold.

Generator (hours run): 14h

Switched generator battery at 5:40pm

Solar— SOC

@ 7:15am : 53%

@ 5:40pm : 98%

Diesel: 50%

Propane: 30%

Ethanol Free Gasoline (5 Gallon containers for ATV): 7.15 Gallons

Water (trailer): 10 Gallons

Water (static): 500 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 129234.7 Gallons

Toilet tank emptied: Yes

ATVs Used: 300, 350.2, 350.3
Oil Added? No
ATV Fuel Used: 1.5 Gallons

# Hours the ATVs were used today: 01:30 hours
Notes on ATVs: ATVs were nominal. We got the 300 working fine.

Deimos rover used: No
Hours: 98.4
Beginning charge: 100%
Ending charge:
Currently charging: Yes

Sojourner rover used: Assigned to director only.
Hours: 5.2
Beginning charge: 100%
Ending charge:
Currently charging: Yes

Spirit rover used: Yes
Hours: 10.5
Beginning charge: 100%
Ending charge: 95%
Currently charging: No

Opportunity rover used: Yes
Hours: 5.5h
Beginning charge: 100%
Ending charge: 62%
Currently charging: No

Curiosity rover used: No
Hours: 3.1
Beginning charge: 100%
Ending charge:
Currently charging: No

HabCar used and why, where? No

General notes and comments: None. Just sitting around in the hab this sol, trying to figure out my VR training module with very limited internet help. It was a cloudy day, so the generator ran all day. We also received a water resupply, yay!

Summary of internet: All nominal
Summary of suits and radios: All nominal
Summary of Hab operations: All nominal
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAM operations: Not Operational
Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: Propane resupply is requested, please.


Melanie Grande, Crew 186 Engineer

Greenhab Report – January 8th

GreenHab Report
Mark Gee

Environmental control:

Shade cloth on

Working Hour: 05:45PM
Inside temp at working hour: 18 C
Outside temp during working hours: 0 C
Inside temperature high: 25 C
Inside temperature low: No Data
Inside humidity: 35 %RH
Inside humidity high: 61 %RH
Inside humidity low: 23 %RH

Hours of supplemental light:
For the crops 05:00 to 11:59 PM

Changes to crops: Kale plants are developing true leaves. The sky has been very cloudy and the plants used very little moisture today.

Daily water usage for crops: 4 gallons

Time(s) of watering for crops: 09:30AM, 04:00PM

Research observations: For the moringa experiment, the plants sprayed with moringa extract seem to be taller. The microgreens are not growing well. I will fertilize them tomorrow.

Changes to research plants: None.

Aquaponics: Not in use.

Narrative: Today I served as the communications officer for the EVA and did not spend much time in the Green Hab.

Support/supplies needed: None.

Science Report – January 8th

Science: Geology

No geology was performed on today’s EVA, except for collection of salts and clay samples in the region of the Moons. However, there was quite a bit of good lab work: thanks to troubleshooting with PANalytical, the TREK portable spectrometer is able again to communicate with the geologist’s laptop which will allow for quantitative analysis of the spectra.
In the meanwhile, work on the samples collected in sol 4 show the richness of the Moons region: Yellow Moon yielded sulfates (mainly gypsum, in the form of selenite), clays (montmorillonite/nontronite, and illite), hematite, and even some light volcanic ash (andesite). Beige Moon have similar composition, with an abundance of large layers of very pure gypsum.
Sulfates and clays are present on Mars, and constitute regions of high geological interest, since they can indicate hydrothermal
paleoenvironment. Human mission could make use of certain types of clays, analogous to those here on Earth, for construction. This makes the research on those materials even more interesting!

EVA Report – January 8th

Author: Max Fagin

Purpose of EVA:

1. Visit the Beige Moon and White Moon areas, which remain unsampled by our previous EVA team during EVA#3

2. Test the Yagi-Uda antenna based navigation experiment by walking back to the hab

Destination: Yellow moon, white moon, beige moon, walk back

UDM27 Coordinates: 516500 E, 4254250 N

Participants: Max Fagin (Commander), Kshitij Mall (Executive Officer), Justin Mansell (Journalist)

Narrative: We exited the hab on schedule and departed north to the Moons region. We started at the western most site, and traveled east, stopping at each site to collect a sample of clay, salt, and a panoramic photo. Despite the cold, not being burdened by sensing equipment was a welcome relief, and we got really good at quickly dismounting, sampling, photographing and documenting a new site. Thanks Cesare for the great geology training! After sampling all three sites, we drove back along Cow Dung road until the GPS indicated we were within 2 km of the hab. We dismounted, set up the radio beacon, and Justin began navigating across the terrain back to the hab, with Kshitij and I following for safety. Inspecting the route on GPS after the fact revealed it was as near to a direct route as could be expected given patches of unnavigable terrain. A great success for the navigation system! All that remains it to duplicate the results with the astronaut’s view of the horizon restricted to remove the possibility of subconsciously navigating by distant terrain features or the sun. We will attempt to do this tomorrow, weather permitting.

After returning to the hab, Justin and Kshitij took three sorties out (one on Opportunity, two on Curiosity) to recover the ATVs that had been abandoned for the navigation experiment and bring them back to the hab.

Note: Another one of our radios gave a low battery signal ~20 minutes into the EVA. Since we were carrying two spare radios as per our decision from yesterday’s EVA, communications were instantly restored. I recommend listing it in the MDRS handbook that the EVA team must always carry at least one fully charged spare radio. It is such a small additional burden, and such a large increase in EVA safety.

Sol Summary – January 8th

Crew 186 Sol 8 Summary Report 08JAN2018

Sol 8

Summary Title: Back to the moon(s of mars)

Author’s name: Max Fagin

Mission Status: All but one of our target EVA sites have been reached, and we are now revisiting sample sites we weren’t able to hit on the first visit. We are also making sure our individual projects will be finished by the end of the mission (just 5 sols from now!).

Sol Activity Summary: Our last EVA to the Moons region was cut short, so we revisited each of the Moons regions and collected a sample of the clay and salt at each site that we had not been able to get the first time. The color contrast from region to region was remarkable to behold; blues, reds, yellows, greys, all due to just small changes in soil chemistry. After we got within 2km of the hab on the way back, we dismounted and tested the radio beacon system to navigate the rest of the way home. Justin and Kshitij then ran a quick three sorties back on the rover to recover the ATVs. Once we were all back at the hab and warmed up, we did some planning for a media event the entire crew will be participating in 2 days from now for students and faculty back at Purdue. Mark spent the day in GreenHab composing some documents to aid future GreenHab officers, Melanie continued the programming of her VR training module, while Sam and Cesare caught up on their own experiments in the science dome.

When we all were back at the hab, Sam filmed a tour of the entire MDRS facility with each of us explaining our roles as part of the crew, then we settled in for an evening episode of ‘Black Mirror’, which has become our crew’s go-to mode for relaxation (And if you’ve ever seen ‘Black Mirror’ you’ll understand why its success at relaxing us is… mixed.)

Look Ahead Plan: Today was the first completely successful test of the navigation system on EVA, we will attempt to replicate the results tomorrow under more challenging conditions on a shorter EVA. The weather forecast is looking less favorable for the next 48 hours, so all EVA plans will be weather contingent.

Anomalies in work: None.

Weather: 1C, winds calm, humidity 80%, Pressure 30.20 inHg

Crew Physical Status: Healthy

EVA: Yellow moon, white moon, beige moon, walk back

Reports to be filed: Sol Summary, Operations Report, Journalist Report, Science Report, Greenhab Report, EVA Report, EVA Request

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site